How many moles are in 3×10233×1023 of aluminium?
1
Expert's answer
2020-11-20T09:59:54-0500

6.023×10^23 aluminum = 1 mole

1 aluminum = (1/6.023×10^23) mole 3×10233×10^23 aluminum

= (3×10233×10^23/6.023×10^23) mole

= 5096.962 mole


So, there are 5096.962 mole of aluminum